OpenStudy (anonymous):

using the substitution method solve 5x+3y=8 y=-2x+1

OpenStudy (anonymous):

So the bottom equation is already solved for y. Plug this into the first equation like so: 5x + 3(-2x+1)=8 Now get rid of the ( ) part: 5x-6x+3=8 Now combine like terms: -x+3=8 Now solve for x: x=-5

OpenStudy (chaise):

5x+3y=8 y=-2x+1 Make y the subject in the first equation. 5x+3y=8 5x-8=-3y (-5/3)x+(8/3)=y Substitute this into your second equation, both equal to y, both equal to eachohter. (-5/3)x+(8/3)=-2x+1 x=-5 Once you do this you must solve for y also. Put your x value into any equation y=-2x+1 y=-2(5)+1 y=-10+1 y=-9

OpenStudy (anonymous):

thanks